S. y=-2+3sin( 21 ) Function 5 Key Point Period: Phase shift Amplitude: Vertical shift: Range:
Answer #1

Ques as y = -2 +3 Sin (2x - 1) y = 3 sin (2x - 2) - 2 ay = 3 Sin (2(x-7) - 2 general form, y = A Sin (B18-c)) +D A=3 Camplitu